[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Aug 24, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Aug 25 02:48:52 UTC 2017


HSU Shift Log for Thu Aug 24 2017 22:48:47 GMT-0400 (EDT)

Shift: R

Forecaster: J Cangialosi

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: After working with Monica, we restored 30 second  
GOES-16 data for Harvey
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170825/2ee7861b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list